GROUPS AS OF WEDNESDAY, SEPT. 9

A ALPHA:  Cloud, Gregg
B BETA:  Fishbeck, Kirkendall, Prinz
Γ GAMMA:  Hamilton, Pitre
Δ DELTA:  Felton, Franz
E EPSILON:  Barkley, Burns
Z ZETA:  Chen, Clearman, Joe
H ETA:  Ruple, Sample
Θ THETA:  Broughton, Mendoza, Skoruppa
I IOTA:  Schaefer, Tidwell
K KAPPA:  Greenlaw, Peck, Portillo
Λ LAMBDA:  Coppinger, Ybarra
M MU:  Lopez, O’Bryant, Pickens
N NU:  Izquierdo, Krueger
Ξ XI:  Cawlfield, Edwards, Marosfalvy
O OMICRON: Qureshi, Southernwood
Π PI:  Billings, Cooper, Dumoit
P RHO:  Randev, ReeseS, Smaistrla
Σ SIGMA:  Bains, Juarez, Romero
T TAU:
Υ UPSILON:  Faasse, Russell
Φ PHI:  Berry, ReeseJ
X CHI:
Ψ PSI:
Ω OMEGA: